Calculate the company's break-even point in sales dollars for the year 2020 if it hires its own sales force to replace the network of agents.Olin Beauty Corporation manufactures cosmetic products that are sold through a network of sales agents. The agents are paid a
Answer #1
Sales = 78015000
Variable costs = 37447200 + (78015000*11%) = 46028850
Fixed costs = fixed COGS + Selling & Marketing exp. + Salesperson FC
7815000 + 10205000 + 5461050 = 23481050
Contribution = Sales - Variable costs = 78015000 - 46028850 = 31986150
Contribution margin %age = contribution / Sales = 31986150/78015000 = 0.41
If the Agents are replaced by Own Salesperson, then
Company's Break-even point in sales dollars = Fixed Costs / Contribution margin
23481050/0.41 = $57270854